# LLaVA-PruMerge: Adaptive Token Reduction for Efficient Large Multimodal Models

### v3

@@ 1 Introduction
Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks. Our work demonstrates the effectiveness of building efficient LMMs from the perspective of visual token reduction and paves the road for further research.
@@ 2.1 Efficient Large Multimodal Models (LMMs)
The need for cross-modal capabilities in resource-limited scenarios has become increasingly important. Despite advancements in LMMs, their large-scale training and deployment incur significant computational costs, necessitating efficient parallel device implementations. Google’s Gemini [Team et al., 2023] is a leader in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices. However, Gemini remains closed-source. Open-source initiatives, like LLaVA-1.5 [Liu et al., 2023a], utilize advanced compression techniques, such as 4/8 bit quantization [Dettmers et al., 2022, Shang et al., 2024]. Further efforts towards efficient LMMs include MobileVLM [Chu et al., 2023], which develops a compact LLM and an efficient multimodal feature projector, and its successor, MobileVLM-v2 [Chu et al., 2024], which explores improved training strategies for mobile scenarios. TinyGPT-V [Yuan et al., 2023a] leverages the advanced Phi-2 [Javaheripi et al., 2023] LLM to surpass the performance of significantly larger models. Similarly, LLaVA-Phi [Zhu et al., 2024] and Vary-toy [Wei et al., 2024] introduce smaller backbones and enhanced vocabularies for broader generalizability. TinyLLaVA [Zhou et al., 2024] investigates the impacts of architectural choices, data quality, and training strategies, demonstrating that smaller LMMs can match the performance of their larger counterparts with optimized data and training. MoE-LLaVA [Lin et al., 2024] adapts Mixture of Experts (MoE) to mitigate model degradation due to sparsity.
@@ 3.1 Preliminaries
Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for large multimodal models, in which the input image is converted into a sequence of representative tokens by the ViT, and then feed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024]. ViTs perform tokenization by dividing an input image into patches and projecting each patch into a token embedding. An extra class token [CLS] is added to the set of image tokens and is responsible for aggregating global image information and performing classification. Let $n$ denote the number of input tokens (image patches) to the ViT. ViTs are composed of a series of transformer blocks. An assembly of several key components composes each transformer block: a multi-head self-attention layer, a feed-forward neural network (FFN), shortcut connections, and layer normalization [Ba et al., 2016], all of which work together to enhance the model’s ability to capture and interpret visual information [Han et al., 2022]. In the self-attention layer, an input token is projected into three distinct vectors: the query vector $\mathbf{q}$, the key vector $\mathbf{k}$, and the value vector $\mathbf{v}$, utilizing three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$. The self-attention mechanism computes the relevance of each item to other items as follows:
@@
where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ represents the dimension of $\mathbf{q},\mathbf{k},\mathbf{v}$. In the last encoder layer of ViT, the [CLS] token is used for classification. Similarly, the attention between [CLS] token and other visual tokens are performed via the attention mechanism:
@@
The multi-head self-attention framework allows for simultaneous focus on multiple positions, offering diverse representation subspaces. This is achieved by employing distinct query, key, and value matrices for different heads, which project the input vectors into varied representation subspaces. Subsequent to the self-attention layers is the feed-forward network (FFN). This network consists of two linear transformation layers separated by a nonlinear activation function and is represented as:
@@
with $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ being the parameter matrices of the linear transformation layers and $\sigma$ denoting the nonlinear activation function, such as GELU [Hendrycks & Gimpel, 2016]. The general forward pass of ViT is illustrated in the left part of Figure 2.
Large Multimodal Models (LMMs). Following the forward pass through a Vision Transformer (ViT), a series of visual tokens is generated. These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which is responsible for mapping the encoded features from other modalities $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$. The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024]. The overall architecture of LMMs is depicted in Figure 1. It is important to note that the computational cost associated with these models tends to increase quadratically with the number of input tokens fed into the LLM backbone [Tay et al., 2022]. There is an increasing demand for the processing of high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs. The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
@@ 3.2 Adaptive Important Token Selection via Outlier Detection
Figure 3: The distribution of attention values (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens. The values in the y-axis are logarithmic values. Notably, most of the spatial visual tokens have near-zero attention values with the class token.
@@
As discussed in Sec. 3.1, LMMs typically leverage an extensive stack of visual tokens to represent the visual information. On the other hand, self-/weakly-supervised learning paradigms, such as CLIP [Radford et al., 2021] simplify this complexity by representing an entire image with a single [cls] token, regarded as the most information-condensed token. To balance those two extreme paradigms, we investigate the Key-Query attention between [cls] token and visual tokens, i.e., $\mathbf{a}_{\texttt{cls}}$ in Equation 3.2. Observing the distribution patterns of attention between the [cls] token and visual tokens unveils a sparse landscape, as depicted in Figure 3. This sparse distribution underpins our methodology for identifying crucial visual tokens. By employing outlier detection algorithms, we aim to adaptively select visual tokens that best represent an image’s features effectively.
@@ 3.3 Token Supplement via Similar Key Clustering
To address this, we devise a token merging method aimed at enhancing the representational capacity of the selected unpruned tokens. This method involves the strategic fusion of currently pruned tokens, as depicted in Figure 2. To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens. Luckily, transformers natively solve this problem with QKV self-attention. Specifically, the keys (K) already summarize the information contained in each token for use in dot product similarity. Thus, we use a dot product similarity metric (e.g., cosine similarity) between the keys of each token to determine which tokens contain similar visual information [Bolya et al., 2023]. Specifically, the similarity between visual tokens is computed as
@@ 3.4 PruMerge+: Bridging the Efficiency-Performance Gap
Figure 4: Image Visualizations of PruMerge and PruMerge+.
PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—without any resultant performance degradation. This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2. A new aspect of PruMerge+ lies in its enhanced token selection process. Beyond merely focusing on the previously identified important tokens, PruMerge+ extends its reach to encompass additional visual tokens from areas initially deemed less critical. This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens. This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 4, thereby minimizing any performance losses while still achieving substantial token reduction.
@@ 4 Experiments
Table 1: Comparison with large multimodal models on six benchmarks. Our proposed approach LLaVA-PruMerge can adaptively reduce visual tokens, which uses only 5.5% visual tokens on average (on 6 tasks) and achieves competitive performance compared to the original LLaVA-1.5.
@@
LLaVA-1.5 + PruMerge+ Vicuna-7B 336 558K 665K 76.5 68.3 57.1 84.0 1431.6 63.5
@@ 4.3.1 Token Sampling Strategy Analysis
Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 5 (d-h). We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (42 tokens), 7 $\times$ 6 (42 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens), .
4.3.2 Effectiveness of Each Module in PruMerge

### v4

@@ 1 Introduction
Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks. Our work demonstrates the effectiveness of building efficient LMMs from the perspective of visual token reduction.
@@ 2.1 Efficient Large Multimodal Models (LMMs)
While LMMs have made significant advances, their large-scale training and deployment incur significant computational costs, requiring efficient parallel device implementations. Google’s Gemini [Team et al., 2023] is a pioneer in efficient LMMs, achieving state-of-the-art performance on multimodal benchmarks and introducing mobile-scale LMMs suitable for low-memory devices, although it is not open-source. Open-source alternatives like LLaVA-1.5 [Liu et al., 2023a] employ advanced compression techniques such as 4/8 bit quantization [Dettmers et al., 2022, Shang et al., 2024]. MobileVLM [Chu et al., 2023] and its improved version, MobileVLM-v2 [Chu et al., 2024], focus on compact architecture designs and training optimizations for mobile use. Similarly, TinyGPT-V [Yuan et al., 2023a] uses the Phi-2 [Javaheripi et al., 2023] LLM backbone to exceed the performance of larger models. LLaVA-Phi [Zhu et al., 2024] and Vary-toy [Wei et al., 2024] introduce smaller backbones and enhanced vocabularies for better generalizability. TinyLLaVA [Zhou et al., 2024] explores the effects of architectural choices and training optimizations, achieving comparable performance to larger models. MoE-LLaVA [Lin et al., 2024] incorporates a Mixture of Experts to address model sparsity, enhancing efficiency and performance.
@@ 3.1 Preliminaries
Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for large multimodal models, in which the input image is converted into a sequence of representative tokens by the ViT, and then feed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024]. An input image is divided into patches and each patch is projected into a token embedding by ViTs. In addition to the set of image tokens, a new class token (i.e., [CLS] token) is added to aggregate global image information and perform classification. Let $n$ denote the number of input tokens (image patches) to the ViT. Let $n$ denote the number of input tokens to the ViT. ViTs consist of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), shortcut connections, and layer normalization [Ba et al., 2016]. These components work together to improve the model’s capability to understand visual data [Han et al., 2022]. In the self-attention layer, an input token is projected into three distinct vectors: query $\mathbf{q}$, key $\mathbf{k}$, and value $\mathbf{v}$, using three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$. The self-attention computes the relevance of each item to other items:
@@
where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ is the dimension of $\mathbf{q}$. In the last layer of ViT, the [CLS] token is used for classification. Similarly, the attention between [CLS] token and other visual tokens is computed by the attention mechanism:
@@
The MSA framework allows for simultaneous attention on multiple positions, offering diverse representation subspaces. This is achieved by employing distinct query, key, and value matrices for different heads, which project the input vectors into different representation subspaces. After the self-attention layers is the feed-forward network (FFN), which consists of two linear transformation layers separated by a nonlinear activation function:
@@
where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the matrices of the linear transformation layers, as well as $\sigma$ denoting the nonlinear activation function. The general forward pass of ViT is illustrated in the left part of Figure 2.
Large Multimodal Models (LMMs). Following the forward pass through a Vision Transformer (ViT), a set of visual tokens is generated. These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which can map the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$. The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024]. The overall architecture of LMMs is depicted in Figure 1. It is important to note that the computational cost with these models increases quadratically with the number of input tokens to the LLM [Tay et al., 2022]. There is an increasing demand for processing of high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs. The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
@@ 3.2 Adaptive Important Token Selection via Outlier Detection
(a) Distribution of attention values.
(b) Image & Tokens Visualizations
Figure 3: (a) The distribution of attention values (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens. The values in the y-axis are logarithmic values. Notably, most of the spatial visual tokens have near-zero attention values with the class token. (b) Image Visualizations of PruMerge and PruMerge+.
@@
As discussed in Sec. 3.1, LMMs typically leverage an extensive stack of visual tokens to represent the visual information. On the other hand, self-/weakly-supervised learning paradigms, such as CLIP [Radford et al., 2021] simplify this complexity by representing an entire image with a single [cls] token, regarded as the most information-condensed token. To balance those two extreme paradigms, we investigate the Key-Query attention between [cls] token and visual tokens, i.e., $\mathbf{a}_{\texttt{cls}}$ in Equation 3.2. Observing the distribution patterns of attention between the [cls] token and visual tokens unveils a sparse landscape, as depicted in Figure 3(a). This sparse distribution underpins our methodology for identifying crucial visual tokens. By employing outlier detection algorithms, we aim to adaptively select visual tokens that best represent an image’s features effectively.
@@ 3.3 Token Supplement via Similar Key Clustering
To address this, we devise a token merging method aimed at enhancing the representational capacity of the selected unpruned tokens. This method involves the strategic fusion of currently pruned tokens, as depicted in Figure 2. To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens. Here we leverage the self-attention mechanism in Vision Transformers (ViT) to metric such a similarity. Specifically, the final layer’s key vector is utilized as the representation of each image token, and thus we use a dot product similarity metric of keys to calculate which tokens have similar visual information [Bolya et al., 2023]. Specifically, the similarity between visual tokens is computed as
@@ 3.4 PruMerge+: Bridging the Efficiency-Performance Gap
PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—without any resultant performance degradation. This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2. A new aspect of PruMerge+ lies in its enhanced token selection process. Beyond merely focusing on the previously identified important tokens, PruMerge+ extends its reach to encompass additional visual tokens from areas initially deemed less critical. This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens. This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing any performance losses while still achieving substantial token reduction.
@@ 4 Experiments
Table 1: Comparison with large multimodal models on six benchmarks. Our proposed approach PruMerge and PruMerge+ can adaptively reduce visual tokens, which use only (respectively) 5.5% and 25.0% visual tokens on average (on 6 tasks) and achieves competitive performance compared to the original LLaVA-1.5.
@@
LLaVA-1.5 + PruMerge+ Vicuna-7B 336 558K 665K 76.8 68.3 57.1 84.0 1462.4 64.9
@@ 4.3.1 Token Sampling Strategy Analysis
Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 4 (d-h). We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 6 $\times$ 7 (42 tokens), 7 $\times$ 6 (42 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).